Spectrum 1460 * TRANSPETTANCE 1120 TLCT D 0821 0862 esti TTTTT 1sbe llwid sample between all plate HAVENUMBERS 99. CopyrightSpectrum 6 19TINOS Cevich 1994 Spectrum F TSMETTANCE 37 BAVENUMERISpectrum G 100 my TRANSMETTINCERI 2000 1500 KRYENUMERII Spectrum H 0.8 TRANSMITTANCE 1000 3000 2000 Wavenumber (cm-1)Spectrum 3000 1000 500 400 2000 1500 Wavenumber(cm-1] Spectrum JIR Molecules Molecule 1 Molecule 2 Molecule 3 Molecule 4 CH CHO Molecule 5 Molecule 6 olololol! OH Molecule 7 Molecule 8 NH2Spectrum A coop NAVENUMERE Spectrum B 188 mom 3215 S - tart BUTT 2 Taal 2920 3382 1882 - 718 T alee sealeases molt betueen sa

Each spectrum corresponds to one of the IR molecules. Match each spectrum to the correct molecule.
Answer #1

1. Benzamide : structure H

  • 1660 cm​​​​​-1 ( C=O stretching frequency)
  • 1400 cm​​​​​-1 ( C-N stretching frequency )
  • 3380 cm​​​-1 ( N-H stretching frequency )

2. pentanal : structure E

  • Around 1700 cm​​​​​-1 (C=O stretching frequency)
  • Around 2750 and 2850 cm​​​​​-1 ( C-H of aldehyde group )
  • Around 2900 cm​​​​​-1 ( C-H stretching frequency of other carbons )

3. Ester : structure C

  • Around 1750 cm​​​-1 ( C=O stretching frequency )
  • Around 1100 to 1200 cm -1 ( C-O stretching frequency )
  • Around 2900 cm​​​​​-1 ( C-H stretching frequency )

4. Phenyl acetlyene : structure G

  • Around 1400 to 1600 cm​​​​​-1 ( C=C for aromatic carbons stretching frequency )
  • Around 3200cm-1 ( C-H of terminal alkyne stretching frequency )
  • Around 2250 cm​​​​​-1 ( C triple bond C stretching frequency)

5. Phenol Ester : Structure A

  • Around 1750 cm​​​-1 (C=O stretching frequency )
  • Around 1100 to 1200 cm​​​​​-1 ( C-O stretching frequency )
  • Around 3300 cm​​​​​-1 ( O-H stretching frequency )

6. Propanoic acid: structure I

  • Around 1710 cm​​​​​-1 ( C=O stretching frequency)
  • Around 1200 cm​​​​​-1 (C-O stretching frequency)
  • Around 2500-3400 cm​​​​​-1 ( O-H stretching frequency )

7. 4methoxyphenol : structure J

  • Around 1400-1600cm-1 ( C=C for aromatic carbon )
  • Around 1100-1200 cm​​​​​-1 ( C-O stretching frequency )
  • Around 3300 cm​-1 (O-H stretching frequency )

8. p-Toulidine : structure B

  • Around 3400cm-1 ( N-H stretching frequency )
  • Around 1400-1600cm-1 ( C=C for aromatic carbon stretching frequency)
  • Around 3000cm-1 ( =C-H stretching frequency)
  • Around 2900cm-1 (-C-H stretching frequency)

9. 4-hydroxy3-methyl2butanone : structure F

  • Around 3300 cm​​​​​-1 ( O-H stretching frequency)
  • Around 1700cm-1 ( C=O stretching frequency )
  • Around 1100-1200 cm​​​​​-1 (C-O stretching frequency)
  • Around 2900cm-1 (-C-H stretching frequency)

10. Cyclohexene : structure D

  • Around 2900cm-1 ( C-H stretching frequency)
  • Around 3000 cm​​​​​-1 ( =C-H stretching frequency)
  • Around 1600cm-1 (C=C stretching frequency)